OpenStudy (anonymous):

Which equation represents the graph of the linear function? A. y = −2x+ 1 B. y = −2x − 1 C. y = 2x + 1 D. y = 2x − 1

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@I_Always_Smiling

OpenStudy (anonymous):

medal and fan plz!!

OpenStudy (anonymous):

if u dont want to give answer plz explain.

OpenStudy (displayerror):

The choices are all in the slope-intercept form \(y=mx+b\) where \(m\) is the slope and \(b\) is the y-intercept. The slope, \(m\), is positive when your line increases going from left to right (i.e. the y-values increase going from left to right). The slope, \(m\), is negative when your line decreases going from left to right (i.e. the y-values decrease going from left to right). The y-intercept, \(b\), is the y-value at which the line crosses the y-axis. Is the slope positive or negative for your line? What is the y-intercept for your line? Knowing these two values will help you answer your question.

OpenStudy (anonymous):

i think its b or d

OpenStudy (displayerror):

Right, the y-intercept is at the point \(\left(0,-1\right)\), so \(b=-1\). What about the slope? Is it positive or negative? Think of the line as the slope of a mountain. If you climb from left to right, are you going up or down? If you are going up, your slope is positive; if you are going down, your slope is negative.

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

?

OpenStudy (displayerror):

Yep! So now you should have your answer.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

if i need more help can i call u out?

OpenStudy (displayerror):

I'll try my best. There are a bunch of people here who can help as well.

OpenStudy (anonymous):

i know but u were the only one that answered.

OpenStudy (displayerror):

You've gotta bump your post so it goes to the top (not sure how though). I had to scroll pretty far down to find your post.

OpenStudy (anonymous):

Complete the equation of the graphed linear function. Write the slope in decimal form. y =?x + -1

OpenStudy (displayerror):

The slope, \(m\), is given by \[\text{Slope} = \frac{rise}{run} =\frac{y_2-y_1}{x_2-x_1}\] Choose two points \(\left(x_1, y_1\right)\) and \(\left(x_2, y_2\right)\), then plug those values into the equation above to get the slope.

OpenStudy (anonymous):

is it 2x-1

OpenStudy (displayerror):

Which two points did you use?

OpenStudy (anonymous):

0,-1 2,0

OpenStudy (displayerror):

Right. So if you plug these two points into the equation I typed above, it should give you \[Slope = \frac{-1 - 0}{0-2}\] Which doesn't equal 2.

OpenStudy (anonymous):

0.5

OpenStudy (displayerror):

Yep! That's your slope.